OpenStudy (anonymous):

Solve the triangle. a = 12, b = 22, C = 95°

OpenStudy (anonymous):

c ≈ 26, A ≈ 27.6°, B ≈ 57.4° c ≈ 26, A ≈ 54.4°, B ≈ 30.6° c ≈ 25.1, A ≈ 27.6°, B ≈ 57.4° c ≈ 25.1, A ≈ 31.6°, B ≈ 53.4°

OpenStudy (anonymous):

law of cosines for this one \[c^2=a^2+b^2-2ab\cos(C)\] \[c^2=12^2+22^2-2\times 12\times 22\cos(95)\] and a calculator

OpenStudy (anonymous):

i get \(c=25.962\) which i guess rounds to 26 http://www.wolframalpha.com/input/?i=c^2%3D12^2%2B22^2-2 \times+12\times+22\cos%2895%29

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

now you can abandon the law of cosines, and move to the law of sines

OpenStudy (anonymous):

\[\frac{\sin(A)}{a}=\frac{\sin(C)}{c}\] \[\frac{\sin(A)}{12}=\frac{\sin(95)}{26}\] \[\sin(A)=\frac{12\sin(95)}{26}\] \[\sin(A)=.46\] \[A=\sin^{-1}(.46)\]

OpenStudy (anonymous):

i get \(27.37\) http://www.wolframalpha.com/input/?i=arcsin%28 \frac{12\sin%2895%29}{26}%29

OpenStudy (anonymous):

yw